Question 1 Kinder Eggs are a favorite chocolate dessert among children. (If you have never had one look them up!) Each Kinder Egg has a surprise inside. The weight of the Kinder Eggs are normally distributed with a mean of weight of 20 grams (g) and a standard deviation of 0.75 grams A. Calculate the probability that a randomly selected kinder egg weight between 18.75 and 21 grams. Show your work. (4 points) You can ship Kinder Eggs to the United States if you purchase a package of 6 eggs. What is the probability that the mean weight of the eggs that you order falls between 18.75 and 21 grams. Show your work. (6 points) B. C. Did you need to know that the population distribution of Kinder Eggs was Normal in order to answer parts (a) or (b)? Justify your answer. (4 points)
Answer #1

Define the standard random variable Z as X-20 0.75 where X denotes the weight of Kinder egg

a) using the normal table we have

18.75- 20 21-20 0.75 P(18.75< X < 21)-P 0.75 P(-1.67 < Z < 1.33) P(Z < 1.33)-P(Z <-1.66) 0.9082-0.0475 0,8607

b) Using central limiting theorem we define the standard random variable Z as

元-20 元-20 0.75//60.31

Hence using the normal table we get

21)-P(1875 1 20 < Z < 20.20 21 -20 0.31 Plǐ8.75 < 0.31-< z < = P(-4.03 < Z < 3.23) = P(Z < 3.23)-P(Z <-4.03) 0.9994- 0.0475 =

c) Yes, population standard deviation is required to define the standard random variable Z.